OpenStudy (anonymous):

Can someone explain on how to solve this? sec^4 2theta?

OpenStudy (anonymous):

what is sec 2 theta?

OpenStudy (anonymous):

@student12345 you have got to answer me now :)

OpenStudy (anonymous):

1 + tan^2 theta

OpenStudy (anonymous):

@surjithayer , @student12345 doesn't know the 2theta's of sin , cos and tan , first ask @student12345 them :) :)

OpenStudy (anonymous):

exactly, that's why i'm having a hard time. i've got to solve this, but no one discussed this to me since i wasn't present.

OpenStudy (anonymous):

\[\sec ^{4}2\theta=\left( \sec ^{2}2\theta \right)^{2}=\left( 1+\tan ^{2}2\theta\right)^{^{2}}\] \[=\left( 1+\left( \frac{ 2\tan \theta }{ 1-\tan ^{2}\theta } \right)^{2} \right)^{2}\] \[=\left( \frac{ \left( 1-\tan ^{2}\theta \right)^{2}+4\tan ^{2}\theta }{ \left( 1- \tan ^{2} \theta\right)^{2} } \right)^{2} \] try further

OpenStudy (anonymous):

\[=\left( \frac{ 1+\tan ^{4}\theta-2\tan ^{2}\theta+4\tan ^{2}\theta }{\left( 1-\tan ^{2}\theta\right)^{2} } \right)^{2}\] I don't know what you want ?

OpenStudy (anonymous):

theta is what im finding.

OpenStudy (anonymous):

What you mean? Give clue.

OpenStudy (anonymous):

either give me answer, so that I can understand.

OpenStudy (anonymous):

i'm finding for theta. what is the angle.

OpenStudy (anonymous):

give me complete statement of question, you are missing something.

OpenStudy (anonymous):

for example, \[2\sec \theta + 1 = \sec \theta\ +3\] \[\sec \theta = 2\] \[\theta = 60°\]

OpenStudy (anonymous):

\[\sec ^{4}2\theta=?\] give me equation

OpenStudy (anonymous):

oh my bad , it's actually \[\sec^42\theta =4\]

OpenStudy (anonymous):

\[\sin 2\theta=2\sin \theta \cos \theta ,\cos 2\theta=\cos ^{2}\theta-\sin ^{2}\theta \] \[Also \sin2\theta =\frac{ 2\tan \theta }{ 1-\tan ^{2}\theta }\]

OpenStudy (anonymous):

i dont get it lol

OpenStudy (anonymous):

\[\sec ^{4}2\theta=4\] \[\frac{ 1 }{ \cos ^{4}2\theta }=4\] \[4\cos ^{4}2\theta=1\] \[\left( 2\cos ^{2} 2\theta \right)^{2}=1\] \[\left( 1+\cos 4 \theta \right)^{2}=1\] \[1+\cos 4 \theta =\pm 1\] \[Either1+\cos 4\theta=1\] \[\cos 4\theta =0 =\cos n \frac{ \pi }{ 2},\theta =\frac{ n \pi }{ 8},where n is an integer\] \[or \cos 4 \theta=-2,regected,\because \left| \cos 4 \theta \right|\le 1\]

OpenStudy (anonymous):

If there is any doubt ask me.